# Swin UNETR++: Advancing Transformer-Based Dense Dose Prediction Towards Fully Automated Radiation Oncology Treatments

### v2

@@ 5 Results
We compared the Swin UNETR’s performance with 3D U-Net (Çiçek et al., 2016), Attention U-Net (Oktay et al., 2018), SegResNet (Myronenko, 2019), DynUNet (a variant of nnU-Net (Isensee et al., 2021)), HD U-net (Nguyen et al., 2019b), C3D (Liu et al., 2021a) and its backboned architecture Swin UNETR (Hatamizadeh et al., 2021). 3D U-Net, Attention U-Net, SegResNet, DynUNet, HD U-net, and C3D are fully CNN-based, among which HD U-net and C3D are specifically designed for 3D radiation dose estimation. The training pipeline for all architectures is the same, except we modified the loss function for C3D to adapt its cascade mechanism. The architecture results measured based on $\overline{S_{\text{DVH}}}$, $\overline{S_{\text{Dose}}}$, $\overline{R_{\text{VA}}}$, and $\overline{R_{\text{PA}}}$ are available in Table 1, where the highest values of each metric in each dataset are highlighted. The details of these results are explained in sections 5.1 and  5.2 from the aspect of ablation study and clinical analysis.
@@ 5.1 Ablation Study
It’s clear from Table 1 that Swin UNETR++ outperforms all seven other architectures on $\overline{S_{\text{DVH}}}$, $\overline{S_{\text{Dose}}}$, $\overline{R_{\text{VA}}}$, and $\overline{R_{\text{PA}}}$ in both validation and test datasets. The only exception is HD U-net’s validation $\overline{R_{\text{VA}}}$ is 2.7% higher than Swin UNETR++’s. Out of 50 more teams, Swin UNETR++ ranks $1^{st}$ on $\overline{S_{\text{DVH}}}$ and $9^{th}$ on $\overline{S_{\text{Dose}}}$ in the validation dataset; ranks $12^{th}$ on $\overline{S_{\text{DVH}}}$ and $10^{th}$ on $\overline{S_{\text{Dose}}}$ in the test dataset.
@@ 5.1.2 The Effect of $\mathcal{L}_{MSE}+\mathcal{L}_{DVH\_global}$
We further examined the advantage of the global DVH loss $\mathcal{L}_{\text{DVH\_global}}$ over the traditional DVH loss $\mathcal{L}_{\text{DVH}}$ described in (Nguyen et al., 2020), which only calculates the structural DVH different based on the PTV and OAR masks, in table 2. These two loss functions are tested on test data since there are more variations in patients’ imaging, thus allowing us to better generalize the effect of $\mathcal{L}_{\text{DVH\_global}}$. $\mathcal{L}_{\text{DVH\_global}}$ demonstrates its efficiency by lowering $\overline{S_{\text{DVH}}}$ by 0.157 Gy and lowering $\overline{S_{\text{Dose}}}$ by 0.159 Gy.
@@ 6 Discussion
This work represents continued efforts to bridge the gap between the technical achievements of computer science and the complexities of clinical medicine. Adopting artificial intelligence in radiation oncology requires clinical endpoints to be at the forefront of collaborations between the different domains. Replacing humans in clinical workflows with artificial intelligence is not appropriate from many patients’ perspectives, and instead, we believe these models will augment the clinical judgment and intelligence of clinical providers. While our results are not state-of-the-art on the specified OpenKBP grand challenge, the difference between the top ten models is very small, and each teams’ training pipeline is significantly different. For example, the No.1 team C3D (Liu et al., 2021a) iterated over each training patient multiple times in each epoch while we only iterated over each training patient once. We also don’t know if some teams used their own patient data to pretrain the model. Based on the experiment we did in section 5, our Swin UNETR has better performance than other models such as C3D (Liu et al., 2021a) and is still representative of a model with clinical viability.
The competition dataset also does not fully represent the data in our clinic, as it contains intensity-modulated radiation therapy (IMRT) plans that use static angles, whereas many clinics have transitioned to the use of volumetric modulated arc therapy (VMAT) plans that represent treatment along continuous arcs. Our clinical data represents more patients, has higher resolutions, contains more organ segmentations, represents our unique patient population, and is more variable in treatment approach than the idealized challenge data. We recognize that this work could benefit from additional clinical testing, but the scope of this project was limited to a summer undergraduate research fellowship. Further clinical training and testing are underway and will provide additional results.

### v3

@@ 5 Results
We compared the Swin UNETR’s performance with 3D U-Net (Çiçek et al., 2016), Attention U-Net (Oktay et al., 2018), SegResNet (Myronenko, 2019), DynUNet, HD U-net (Nguyen et al., 2019b), C3D (Liu et al., 2021a) and its backboned architecture Swin UNETR (Hatamizadeh et al., 2021). 3D U-Net, Attention U-Net, SegResNet, DynUNet, HD U-net, and C3D are fully CNN-based, among which HD U-net and C3D are specifically designed for 3D radiation dose estimation. The training pipeline for all architectures is the same, except we modified the loss function for C3D to adapt its cascade mechanism. The architecture results measured based on $\overline{S_{\text{DVH}}}$, $\overline{S_{\text{Dose}}}$, $\overline{R_{\text{VA}}}$, and $\overline{R_{\text{PA}}}$ are available in Table 1, where the highest values of each metric in each dataset are highlighted. The details of these results are explained in sections 5.1 and  5.2 from the aspect of ablation study and clinical analysis.
@@ 5.1 Ablation Study
It’s clear from Table 1 that Swin UNETR++ outperforms all seven other architectures on $\overline{S_{\text{DVH}}}$, $\overline{S_{\text{Dose}}}$, $\overline{R_{\text{VA}}}$, and $\overline{R_{\text{PA}}}$ in both validation and test datasets. The only exception is HD U-net’s validation $\overline{R_{\text{VA}}}$ is 2.7% higher than Swin UNETR++’s. Out of 50 more teams, Swin UNETR++ should rank $1^{st}$ on $\overline{S_{\text{DVH}}}$ and $9^{th}$ on $\overline{S_{\text{Dose}}}$ in the validation dataset; ranks $12^{th}$ on $\overline{S_{\text{DVH}}}$ and $10^{th}$ on $\overline{S_{\text{Dose}}}$ in the test dataset on the leaderboard.
@@ 5.1.2 The Effect of $\mathcal{L}_{MSE}+\mathcal{L}_{DVH\_global}$
We further examined the advantage of the global DVH loss $\mathcal{L}_{\text{DVH\_global}}$ over the traditional DVH loss $\mathcal{L}_{\text{DVH}}$, which only calculates the structural DVH different based on the PTV and OAR masks, in table 2. These two loss functions are examined on test data since there are more variations in patients’ imaging, thus allowing us to better generalize the effect of $\mathcal{L}_{\text{DVH\_global}}$. $\mathcal{L}_{\text{DVH\_global}}$ demonstrates its efficiency by lowering $\overline{S_{\text{DVH}}}$ by 0.157 Gy and lowering $\overline{S_{\text{Dose}}}$ by 0.159 Gy.
@@ 6 Discussion
This work represents continued efforts to bridge the gap between the technical achievements of computer science and the complexities of clinical medicine. Adopting AI in radiation oncology requires clinical endpoints to be at the forefront of collaborations between the different domains. Replacing humans in clinical workflows with artificial intelligence is not appropriate from many patients’ perspectives, and instead, we believe these models will augment the clinical judgment and intelligence of clinical providers. While our results are not state-of-the-art on the specified OpenKBP grand challenge, the difference between the top ten models is very small, and each teams’ training pipeline is significantly different. For example, the No.1 team C3D iterated over each training patient multiple times in each epoch while we only iterated over each training patient once. We also don’t know if some teams used their own patient data to pretrain the model. Based on the experiment we did in section 5, our Swin UNETR has better performance than other models such as C3D (Liu et al., 2021a) and is still representative of a model with clinical viability.
The competition dataset also does not fully represent the data in our clinic, as it contains IMRT plans that use static angles, whereas many clinics have transitioned to the use of volumetric modulated arc therapy (VMAT) plans that represent treatment along continuous arcs. Our clinical data represents more patients, has higher resolutions, contains more organ segmentations, represents our unique patient population, and is more variable in treatment approach than the idealized challenge data. We recognize that this work could benefit from additional clinical testing, but the scope of this project was limited to a summer undergraduate research fellowship. Further clinical training and testing are underway and will provide additional results.